25 April (Monday)

  • Acitvity
    • Reco16 validation production on Full stream first data launched

  • Site Issues
    • T0:
    • T1:
      • SARA transfer problems (GGUS:121005) fixed
      • GRIDKA SRM problem (GGUS:120968) and observed very low staging throughput

18 April (Monday)

  • Acitvity
    • Staging activities on all T1 sites for 2012 data stripping
  • Site Issues
    • T0: Problem uploading to EOS Disk , fixed on 12th (GGUS:120799)
    • T1:
      • SARA FTS problems when staging in data (GGUS:120808), fixed
      • IN2P3 gridftp problem to access files (GGUS:120835), fixed
      • RAL network problem is fixed
      • RAL disk server is still being investigated (some files inaccessible)
      • GRIDKA running few jobs is fixed (increased priority for the jobs) (GGUS:120874), fixed
      • GRIDKA: GGUS was unstable this morning, back already

11 April (Monday)

  • Site Issues
    • T0: Problem with files accessing EOS Disk, solved during the week-end
    • T1:

4 April (Monday)

  • Activity
    • Very high activity
      • Incremental Run1 Stripping Campaign in full swing
      • Turbo data reprocessing half done
      • MC and User jobs

  • Site Issues
    • T1:
      • FZK-LCG2: Problem uploading job output data to local storage (GGUS:120533)
      • INFN-T1: Problem uploading job output data to local storage (GGUS:120530)

-- JoelClosier - 2017-01-10

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2017-01-10 - JoelClosier
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback